Two systems have the following equations of state, where R is the gas constant per mole. The mole numbers are N1 = 3 and N2 = 5. The initial temperatures are T1 = 175 K and T2 = 400 K. What is the temperature once equilibrium is reached (in Kelvin)?
1
Expert's answer
2020-07-27T07:23:39-0400

we know that, for two particular system we have,"T_1=175 K,T_2=400K"

"n_1=3,n_2=5,"

we know that

"\frac{1}{T}=\frac{3}{2}R (\frac{n}{U})"


for two different temperature

"\frac{1}{T_1}=\frac{3}{2}R (\frac{n_1}{U_1})"


"\frac{1}{175}=\frac{3}{2}R (\frac{3}{U_1})"


"U_1=6547.275"

similarly for other system

"\frac{1}{T_2}=\frac{3}{2}R (\frac{n_2}{U_2})"


"\frac{1}{400}=\frac{3}{2}R (\frac{5}{U_2})"


"U_2=24942 J"

so,

"U=U_2-U_1=24942-6547.275"

so, during equilibrium condition

"\frac{1}{T}=\frac{3}{2}R (\frac{n}{U})"


"\frac{1}{T}=\frac{3}{2}(8.314)(\frac{8}{18394.725})"


T= 184.375 K

This is the equilibrium temperature